OpenStudy (anonymous):

Find the solution to the equation: -4 (y-5) = 5 ( y - 4)

OpenStudy (anonymous):

-4y+20=5y-20 -4y-5y=-20-20 -9y=-40 y=40/9

OpenStudy (anonymous):

is this helpful?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Thanks